OpenStudy (anonymous):

Sarah’s dad tells her she can have the greater of the following two amounts for her allowance (in dollars): The sum of 3 times 5 and 2 3 times the sum of 5 and 2 What is Sarah’s allowance, in dollars?

OpenStudy (anonymous):

sum of 3 times 5 and 2 is 3*5 +2 = 17 3times the sum of 5 and 2 = 3(5+2) = 3*7 = 21 2nd one